OpenStudy (anonymous):

A spherical balloon is being inflated. Find the volume of the balloon at the instant when the rate of increase of the surface area is 8 times the rate of increase of the radius of the sphere.

OpenStudy (anonymous):

\[\S(r)=4\pi*r^2\] \[V(r)=(4/3)\pi*r^3\] Implicitly differentiate the S(r) w.r.t. t, time, to compute the rate of increase of surface area and radius. \[dS/dt=(dr/dt)*8\pi*r\] \[dS/dt=8*dr/dt\] \[8*dr/dt=(dr/dt)*8\pi*r\] \[r=1/\pi\] \[V=(4/3)*\pi*(1/\pi)^3\] Volume is approximately equal to .135095 Not 100% sure on this.